OpenStudy (anonymous):

Find the general solution to the differential equation. Prime denots derivative with respect to x. (x+y)y'=x-y The answer is x^2-2yx-y^2 I'm a little confused as to how to sub. I'm pretty sure that this uses Beroulli equation.

OpenStudy (experimentx):

write y' as dy/dx .. and bring both to same side and see if you see something.

OpenStudy (anonymous):

I get \[\frac{ dx }{ dy } = \frac{ x }{ x+y } - \frac{ y }{ x+y }\] I should move the y/x+y to the other side andmove the dx and dy?

OpenStudy (anonymous):

*dy/dx

OpenStudy (experimentx):

 does this ring the bell?